OpenStudy (anonymous):

Simplifying Radicals 1. 2√125 2. 5√60 3. √y^4 4. √x^2y 5. √50a^8 6. √12a^3b^6

OpenStudy (anonymous):

factorize the number inside the radical

OpenStudy (anonymous):

how do i do that? find the square root? @idvez

OpenStudy (anonymous):

lets take the first 125, what are the factors of 125?

OpenStudy (anonymous):

11.18033989 @idvez

OpenStudy (anonymous):

factor of a number is when you multiply two other numbers to give you the initial one for example 4 factors into 2 and 2, because 2*2 =4

OpenStudy (anonymous):

so what two numbers can you multiply to give you 125

OpenStudy (anonymous):

1,5,25 @idvez

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so lets rewrite 125 under the radical as \[2\sqrt{5*25*1}=2\sqrt{5*25}\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

what is confusing you?

OpenStudy (anonymous):

like the whole factoring in general

OpenStudy (anonymous):

i took your first problem \[2\sqrt{125}\] and replace 125 with the factor you gave me, ant the end i did not writhe 1 because any number multiplied by one is the same number

OpenStudy (anonymous):

ok... lets take another number 36, what are the factor of 36

OpenStudy (anonymous):

i like to start always with one, what number do i have to multiply 1 by to get 36?

OpenStudy (anonymous):

for future reference factors of 36 1 and 36 2 and 18 3 and 12 4 and 9 6 and 6 factors:1,2,3,4,6,36,18,12,9

OpenStudy (anonymous):

get very conformable breaking a number into its factor as this will help you tremendously for the kind of problems you were asking...

OpenStudy (anonymous):

the next step after you factorize the number under the radical,is to see if you can take any factors out of the radical, you can do that by further factorizing into prime factors or by knowing some simple square roots such as sqrt(25)=5